OpenStudy (anonymous):

Which exponential function goes through the points (1, 32) and (4, 2048)? f(x) = 4(4)x f(x) = 8(2)-x f(x) = 8(4)x f(x) = 4(4)-x

OpenStudy (amistre64):

none of those appear to be exponential functions ....

OpenStudy (anonymous):

f(x) = 4(4)^x f(x) = 8(2)^-x f(x) = 8(4)^x f(x) = 4(4)^-x

OpenStudy (amistre64):

better :) now tell me, what do we know about a negative exponent, do they get bigger in value or smaller in value for x>1?

OpenStudy (anonymous):

Smaller?

OpenStudy (amistre64):

smaller is correct, even tho i worded the question almost badly; the options given would get smaller for negative exponents ... since the value at x=4 is bigger than at x=1, we can rule out 2 options, now test the other 2 for when x=1

OpenStudy (anonymous):

It's not the last one either right?

OpenStudy (anonymous):

Is it A?

OpenStudy (amistre64):

yeah, 2nd and 4th ones are just plan off when x=1, the other 2 options are simple enough to determine

OpenStudy (amistre64):

4(4)^1 = 16, not 32 ... so lets avoid A

OpenStudy (anonymous):

So C would be the best option? I'm really bad at Math.

OpenStudy (amistre64):

:) C is what I would pick yes: 8(4)^1 = 32 ... and we want (1,32) so it fits the best

OpenStudy (anonymous):

thanks!

OpenStudy (amistre64):

yw